OpenStudy (anonymous):

The sum of two consecutive odd integers equals 12. What are they? x represents integer x + 2 represents integer

OpenStudy (mrnood):

so one of the numbers is x, the other is x+2 (because they are consecutive ODD numbers) so write down an expression that is the sum of those 2 numbers Write oyur answer here and I will help with the next step

OpenStudy (anonymous):

\[x+2=?\]

OpenStudy (anonymous):

@CausticSyndicalist @iGreen @MrNood

OpenStudy (anonymous):

Anyone, plz? (:

OpenStudy (mrnood):

x+ 2 is ONE of the numbers now write down that number PLUS the other one

OpenStudy (mrnood):

@xo.minnie.xox soz - that is not correct

OpenStudy (anonymous):

Okay lol. How do you do it?

OpenStudy (mrnood):

If @messenger.Nat<3 will do the first step (write down the expression for adding the two numbers) then I will show the next step

OpenStudy (anonymous):

The sum of two consecutive odd integers equals 12. What are they? x represents the a) larger integer or b) smaller integer x + 2 represents the a) larger integer or b) smaller integer

OpenStudy (mrnood):

If you want help with the first bit - then do what I asked before ONE number is x another is (x+2) Write down the expression for ADDING the two numbers together

OpenStudy (anonymous):

@MrNood I don't understand what you mean... in Algebra, the lesson is on WRITING EQUATIONS FROM WORD PROBLEMS- Algebra 1

OpenStudy (mrnood):

Exactly - so I am trying to get you to do just htat. The question tells you that 2 numbers add up to 12 we don't know what those numbers are so in algebra we write a letter for unknowns so we could write x+ y = 12 BUT we know more than that - because it tells you that the 2 numbers are consecutive ODD numbers so if one of the numbers is x then we know the other must be x+2 so now you can write x+y = 12 in a different way because we know that the other number that we originally called y can alternativelt be written as x+2 SO write x+y=12 but put in the other value that we now have for y... (x+2)

OpenStudy (mrnood):

IF you just want me to give you the answer then say so now...

OpenStudy (anonymous):

@MrNood nah, it's okay, I figured it out after reading the question for the 15th time xD Thanks, though!

OpenStudy (mrnood):

That's good (I wasn't going to give the answer anyhow - just wanted to know if that was what you were waiting for..)